# Quantum Networks: from Multipartite Entanglement to Hypergraph Immersion

### v1

@@ Appendix A Supplemental Materials
We consider the important notion of immersion in this paper. Specifically, for ordinary graphs, an immersion of $H$ in $G$ is a function $\alpha$ with domain $V(H)\cup E(H)$ such that: (i) $\alpha(v)\in V(G)$ for all $v\in V(H)$, and $\alpha(u)\neq\alpha(v)$ for all distinct $u,v\in V(H)$; (ii) for each edge $e\in E(H)$, if $e$ has distinct ends $u,v$ then $\alpha(e)$ is a path of $G$ with ends $\alpha(u),\alpha(v)$; (iii) for all distinct $e,f\in E(H)$, $E(\alpha(e)\cap\alpha(f))=\emptyset$.
The notion of embedding (a.k.a. topological minor) is closely related to immersion. Specifically, an embedding of $H$ in $G$ is a function $\beta$ that is defined the same as an immersion $\alpha$, except that in rule (iii), $E(\beta(e_{1})\cap\beta(e_{2}))=\emptyset$ is replaced by $V(\beta(e_{1})\cap\beta(e_{2}))=\beta(V(e_{1}\cap e_{2}))$ where $V(e_{1}\cap e_{2})$ denotes the common endpoints of $e_{1}$ and $e_{2}$. In other words, $\beta$ is an injective mapping from edges in $H$ to vertex-disjoint paths, instead of edge-disjoint paths as in immersion. As a special example, it is well known that an embedding of the complete graph of size $4$ exists if and on if the graph is not series-parallel. We note that an embedding is an immersion, but not vice versa.
@@
The connection in hypergraphs can be considered through the generalisation of a path in an ordinary graph to hypergraphs, the Berge path [52]. A Berge path of length $t$ is an alternating sequence of distinct $t+1$ vertices and distinct $t$ hyperedges of the hypergraph $G$, $v_{1},e_{1},v_{2},e_{2},v_{3},\cdots,e_{t},v_{t+1}$, such that $v_{i},v_{i+1}\in e_{i}$, for $i=1,\cdots,t$. A Berge cycle is the same as a Berge path except that the last vertex $v_{t+1}$ is replaced by $v_{0}$, making the sequence cyclic. A connected subgraph of a hypergraph is then a set of vertices which are pairwise connected by some Berge path(s). The size of the subgraph is the size of its vertex set.
Appendix C 2. Criteria on Immersion of $K_{4}^{2}$ and $K_{4}^{3}$
In this section, we discuss in more detail how to determine whether there is an immersion of $K_{4}^{2}$ and $K_{4}^{3}$.
C.1 $K_{4}^{2}$
To determine the immersion of $K_{4}^{2}$, we utilise the fact that $K_{4}^{2}$ is an ordinary graph. Specifically, we consider the strategy to apply a series of evaporation operations to the edges of size greater than $2$ in the hypergraph of interest, $G$, until it becomes an ordinary graph (where every edges have size $2$), $\bar{G}$. We denote the set of all such ordinary graphs obtained from hypergraph $G$ by $\mathcal{S}=\{\bar{G}_{i}\}_{i=1}^{k}$. It is straightforward to show that an immersion of $K_{4}^{2}$ exists in $G$ (as defined in hypergraphs) exists if and only if an immersion of $K_{4}^{2}$ exists in $\bar{G}_{i}$ for some $i$ (as defined in ordinary graphs). Therefore, if there is some $i$ such that $\bar{G}_{i}$ is non-series-parallel (see [53] for an efficient algorithm), then an embedding of $K_{4}^{2}$ exists, thus an immersion of $K_{4}^{2}$ exists; Else, for every three-edge-connected components of each $\bar{G}_{i}$, run Algorithm 1 to check whether any of them fall into the special cases where an immersion of $K_{4}^{2}$ exists (see [46] for more details on constructing three-edge-connected components). Specifically, we will apply the pruning operation: in a three-edge-connected series-parallel multigraph, suppose $v$ is a vertex with exactly two neighbours $u$ and $w$, and suppose there is only one copy of edge $uw$ (then at least two copies of $uv$ by three-edge-connectivity); we say that $v$ is pruned if the multiplicity of $uv$ is set to $2$, and we say that a graph is pruned if each vertex fitting the profile of $v$ is pruned.
Algorithm 1 test($\bar{G}$).
1:Input: a three-edge-connected series parallel multigraph $\bar{G}$.
2:Output: YES, if $\bar{G}$ contains an immersed $K_{4}$, NO otherwise.
3:for each vertex $v$ in $\bar{G}$ with exactly one neighbour do
@@ C.1 $K_{4}^{2}$
6:if any cut point in $\bar{G}$ has degree $7$ or more then
@@
C.2 $K_{4}^{3}$
We first note that any edges of higher orders, i.e., consisting of more than $3$ nodes, contain redundant information for the immersion of $K_{4}^{3}$ where all edges have size $3$. Hence, in the following, we focus on hypergraphs where hyperedges have at most $3$ nodes. In particular, we take a first step to list all possible variants of hypergraphs of size $4$ with an immersion of $K_{4}^{3}$, that are minimal, in the sense that no edges in the variants remain unused for defining the immersion; see Fig. 3. A hypergraph of interest, $G$, has an immersion of $K_{4}^{3}$ if and only if $G$ has an immersion of any of the $18$ variants in Fig. 3.
Figure 3: The minimum subset $\tilde{\mathcal{D}}(H)$. For $H=K_{4}^{3}$, the minimum subset $\tilde{\mathcal{D}}(K_{4}^{3})$ includes $18$ variants of different topologies. Including all variants ensures that, for every $G$ in which $H$ has an immersion, there is at least one $\tilde{H}\in\tilde{\mathcal{D}}(H)$ such that $\tilde{H}$ has an immersion in $G$.

### v2

@@ Abstract
We consider the important notion of immersion in this paper. Specifically, for ordinary graphs, an immersion of $H$ in $G$ is a function $\alpha$ with domain $V(H)\cup E(H)$ such that [29]:
  1. 1.
    $\alpha(v)\in V(G)$ for all $v\in V(H)$, and $\alpha(u)\neq\alpha(v)$ for all distinct $u,v\in V(H)$;
  2. 2.
    for each edge $e\in E(H)$, if $e$ has distinct ends $u,v$ then $\alpha(e)$ is a path of $G$ with ends $\alpha(u),\alpha(v)$;
  3. 3.
Equivalently, an immersion of $H$ in $G$ exists if and only if $H$ can be obtained from (a subgraph of) $G$ by lifting, an operation on adjacent edges: given three vertices $v$, $u$, and $w$, where $\{v,u\}$ and $\{u,w\}$ are edges in the graph, the lifting of $v,u,w$, or equivalently of $\{v,u\},\{u,w\}$ is the operation that deletes the two edges $\{v,u\}$ and $\{u,w\}$ and adds the edge $\{v,w\}$.
The notion of embedding (a.k.a. topological minor) is closely related to immersion. Specifically, an embedding of $H$ in $G$ is a function $\beta$ that is defined the same as an immersion $\alpha$, except that in rule (iii), $E(\beta(e_{1})\cap\beta(e_{2}))=\emptyset$ is replaced by $V(\beta(e_{1})\cap\beta(e_{2}))=\beta(V(e_{1}\cap e_{2}))$ where $V(e_{1}\cap e_{2})$ denotes the common endpoints of $e_{1}$ and $e_{2}$. In other words, $\beta$ is an injective mapping from edges in $H$ to vertex-disjoint paths, instead of edge-disjoint paths as in immersion. As a special example, it is well known that an embedding of a complete graph of size $4$ exists if and on if the graph is not series-parallel [47]. We note that an embedding is an immersion, but not vice versa.
@@
The end-to-end connection in hypergraphs can be considered through the generalization of a path in an ordinary graph to hypergraphs, the Berge path [46]. A Berge path of length $t$ is an alternating sequence of distinct $t+1$ vertices and distinct $t$ hyperedges of the hypergraph $G$, $v_{1},e_{1},v_{2},e_{2},v_{3},\cdots,e_{t},v_{t+1}$, such that $v_{i},v_{i+1}\in e_{i}$, for $i=1,\cdots,t$. A Berge cycle is the same as a Berge path except that the last vertex $v_{t+1}$ is replaced by $v_{0}$, making the sequence cyclic. A connected subgraph of a hypergraph is then a set of vertices which are pairwise connected by some Berge path(s). The size of the subgraph is the size of its vertex set.
Section II Criteria on Immersion of $K_{4}^{2}$ or $K_{4}^{3}$
This section details how to determine whether $K_{4}^{2}$ or $K_{4}^{3}$ can be immersed in a hypergraph $G$. Our main strategy is to transform the hypergraph immersion problem for $G$ into a set of (simpler) immersion problems on a new set of ordinary graphs, denoted $\{G^{\text{dw}}_{i}\}_{i}$ (Fig. S1). Each $G^{\text{dw}}_{i}$ is uniquely derived from $G$, containing only size-$2$ edges. This reduction simplifies the problem, as the immersion of the edges in $G^{\text{dw}}_{i}$ can then be treated using the lifting operations, opening up the possibility of leveraging existing techniques and theorems.
While the following criteria—as we will present below—are intuitive, they do not readily generalize to hypergraphs beyond $K_{4}^{2}$ and $K_{4}^{3}$. Consequently, a general reduction of hypergraph immersion to ordinary graph immersion is not straightforward. Moreover, the associated testing procedures may not be as efficient and can exhibit exponential time complexity w.r.t. the size of $G$. A proof demonstrating the existence of a more efficient, polynomial-time algorithm for general hypergraph immersion can be found in Ref.  [^2] In preparation..
Figure S1: Conversion of a hypergraph $G$ into sets of ordinary graphs $\{G^{\text{dw}}_{i}\}_{i}$. Two example graphs are shown, one from the set $\{G^{\text{dw:}\{2\}}_{i}\}_{i}$ and one from $\{G^{\text{dw:}\{2,3^{\prime}\}}_{i}\}_{i}$.
A $K_{4}^{2}$
To determine if $H=K_{4}^{2}$ can be immersed in $G$, we consider the set $\{G^{\text{dw:}\{2\}}_{i}\}_{i}$ of all possible ordinary graphs derivable from $G$ solely through dewetting operations (Fig. S1). Each graph $G^{\text{dw:}\{2\}}_{i}$ retains the original number of hyperedges (i.e., $\left|E(G^{\text{dw:}\{2\}}_{i})\right|=\left|E(G)\right|$), but after dewetting, each hyperedge $e\in E(G)$ loses connection to all but only two incident nodes, becoming an ordinary edge. The following theorem provides a criterion for the immersion of $K_{4}^{2}$:
Theorem II.1.
The complete ($2$-uniform) graph $H=K_{4}^{2}$ can be immersed in a hypergraph $G$ if and only if there exists a subgraph of $G^{\text{dw:}\{2\}}_{i}$ for some $i$, such that $K_{4}^{2}$ can be constructed by lifting operations applied to the edges of the subgraph of $G^{\text{dw:}\{2\}}_{i}$.
Proof.
The sufficient condition (“if”) is straightforward. Since the lifting operation is nothing but a combination of the coalescence and dewetting operations for ordinary graphs, it implies that $H=K_{4}^{2}$ is derivable by coalescence and dewetting operations on a subgraph of $G^{\text{dw:}\{2\}}_{i}$, which is derivable by dewetting operations on $G$. This, by definition, constitutes an immersion of $K_{4}^{2}$ in $G$.
The necessary condition (“only if”) goes as follows. Given an immersion $\alpha$ of $H$ in $G$, for each edge $e=\{u,v\}\in H=K_{4}^{2}$, the corresponding connected subgraph $\alpha(e)$ in $G$ must contain a Berge path that connects $u$ and $v$, which can be derived from $\alpha(e)$ through dewetting. This Berge path, composed entirely of size-2 edges, can be lifted to form the corresponding edge $e\in K_{4}^{2}$. Since for all distinct $e_{1},e_{2}\in E(H)$, $\alpha(e_{1})$ and $\alpha(e_{2})$ are edge-disjoint, the corresponding Berge paths are also edge-disjoint. Therefore, the union of the Berge paths of $\alpha(e)$ for every $e\in E(H)$ must be a subgraph of some member of the set $\{G^{\text{dw:}\{2\}}_{i}\}_{i}$. ∎
Consequently, the hypergraph immersion problem for $G$ is reduced to a set of ordinary graph immersion problems, one for each $G^{\text{dw:}\{2\}}_{i}$. The immersion of $H=K_{4}^{2}$ in ordinary graphs is a well-studied problem [48]. Specifically, if $G^{\text{dw:}\{2\}}_{i}$ is non-series-parallel (see Ref. [55] for an efficient algorithm), then an embedding of $K_{4}^{2}$ exists, and thus an immersion of $K_{4}^{2}$ exists; Else, for every three-edge-connected components of each $G^{\text{dw:}\{2\}}_{i}$, run Algorithm 1 to check whether any of them falls into the special cases where an immersion of $K_{4}^{2}$ exists (see Ref. [48] for more details on constructing three-edge-connected components).
Algorithm 1 test($X$).
1:Input: a three-edge-connected series-parallel graph $X$.
2:Output: YES, if $X$ contains an immersed $K_{4}$, NO otherwise.
3:for each vertex $v$ in $X$ with exactly one neighbour do
@@ A $K_{4}^{2}$
6:if any cut point in $X$ has degree $7$ or more then
@@
In particular, the pruning operation in Algorithm 1 is defined as follows: in a three-edge-connected series-parallel multigraph, suppose $v$ is a vertex with exactly two neighbours $u$ and $w$, and suppose there is only one copy of edge $\{u,w\}$ (then at least two copies of $\{u,v\}$ by three-edge-connectivity); we say that $v$ is pruned if the multiplicity of $\{u,v\}$ is set to $2$, and we say that a graph is pruned if each vertex fitting the profile of $v$ is pruned.
Figure S2: Variants of $H=K_{4}^{3}$. The set $\mathcal{F}(H)$ contains $18$ distinct graph topologies derived from $K_{4}^{3}$. Each size-$3$ hyperedge in $K_{4}^{3}$ is replaced by either two edges connecting the three incident vertices or a Y-shaped graph with an additional vertex.
B $K_{4}^{3}$
Similarly, to determine if $H=K_{4}^{3}$ can be immersed in $G$, we consider the set of all possible hypergraphs derivable from $G$ solely through dewetting operations, such that each hyperedge $e\in E(G)$ reduces to either a size-$2$ edge or a size-$3$ hyperedge. Then, we replace each size-$3$ hyperedge $e$ by its factor graph $e^{\prime}$, i.e., a Y-shaped graph with an additional central vertex denoted by “$\times$” (Fig. S1). The set of all possible graphs derived this way is denoted by $\{G^{\text{dw:}\{2,3^{\prime}\}}_{i}\}_{i}$.
In addition, we construct a set of ordinary graphs, $\mathcal{F}(H)$, derived from $H=K_{4}^{3}$ by replacing each size-$3$ hyperedge with either two ordinary edges connecting the three incident vertices or a Y-shaped graph (three edges connecting the vertices to an additional central vertex). Up to isomorphism, this yields $18$ distinct graph topologies, illustrated in Fig. S2. It is straightforward to see that for every $F\in\mathcal{F}(H)$, there exists an immersion of $H$ in $F$.
The following theorem provides a criterion for the immersion of $K_{4}^{3}$:
Theorem II.2.
The complete $3$-uniform hypergraph $H=K_{4}^{3}$ can be immersed in a hypergraph $G$ if and only if there exists a subgraph of $G^{\text{dw:}\{2,3^{\prime}\}}_{i}$ for some $i$, such that at least one of the $18$ graphs $F\in\mathcal{F}(H)$ (Fig. S2) can be derived from the subgraph of $G^{\text{dw:}\{2,3^{\prime}\}}_{i}$ by lifting operations, with the restriction that no vertex in $H$ will be mapped to a vertex in $G^{\text{dw:}\{2,3^{\prime}\}}_{i}$ that has a “$\times$” label.
Proof.
Similar to Theorem II.1, the sufficient condition (“if”) is as follows. If one of the graphs $F\in\mathcal{F}(H)$ can be immersed in $G^{\text{dw:}\{2,3^{\prime}\}}_{i}$, then by construction, $H$ can also be immersed in $G^{\text{dw:}\{2,3^{\prime}\}}_{i}$. We must now verify that this immersion remains valid when extended to $G$. Since “$\times$” vertices are not actual vertices in $G$, two potential issues could invalidate the immersion: (1) a vertex of $H=K_{4}^{3}$ is mapped to a “$\times$” vertex, or (2) a lifting operation involving a “$\times$” vertex creates a topology not derivable from $G$ through coalescence and dewetting. The first issue is precluded by the theorem’s conditions. The second is also impossible: at most three edge-disjoint paths can terminate at a “$\times$” vertex. Lifting joins two of these paths, while the third becomes isolated and irrelevant to $H$’s immersion (since “$\times$” vertices cannot map to vertices in $H$). This isolated path can thus be discarded. The joined two paths, however, are equivalent to dewetting the original size-$3$ hyperedge in $G$ to a size-$2$ edge, followed by lifting operations. Therefore, the immersion of $H$ in $G^{\text{dw:}\{2,3^{\prime}\}}_{i}$ (with the “$\times$” vertex preclusion) implies sufficient connectivity for $H$ to be immersed in $G$.
The necessary condition (“only if”) is as follows. For each size-$3$ hyperedge $e=\{u,v,w\}\in K_{4}^{3}$, the corresponding connected subgraph $\alpha(e)$ in $G$ contains three Berge paths (from $u$ to $v$, $v$ to $w$, and $w$ to $u$, respectively). There are only three possibilities for these paths: (1) Two Berge paths are edge-disjoint (e.g., $u$ to $v$ and $v$ to $w$ w.l.o.g.). (2) No two Berge paths are edge-disjoint, and an additional vertex $x$ exists in $\alpha(e)$ with edge-disjoint sub-paths connecting $x$ to $u$, $x$ to $v$, and $x$ to $w$, respectively. (3) No two Berge paths are edge-disjoint, and at least one hyperedge in $\alpha(e)$ is shared by all three Berge paths.
In case (1), the two Berge paths can be derived from $\alpha(e)$ by dewetting to size-$2$ edges. In case (2), similarly, the three sub-paths can be derived by dewetting to all size-$2$ edges. In case (3), we dewet $\alpha(e)$ to all size-$2$ edges, except for one of the shared hyperedges. This remaining hyperedge is then dewetted to a size-$3$ hyperedge connecting exactly to $u$, $v$, and $w$ via edge-disjoint sub-paths. Replacing this size-3 hyperedge with its factor graph makes this case equivalent to the second case, with the central vertex corresponding to $x$. Since in all three cases the Berge paths (or sub-paths involving $x$) are edge-disjoint, the union of them must be a subgraph of some member of the set $G^{\text{dw:}\{2,3^{\prime}\}}_{i}$.
Finally, the construction of $\mathcal{F}(H)$ enumerates all possible combinations of these cases (two-edge-series graph or Y-shaped graph) for each hyperedge, up to isomorphism. This completes our proof. ∎
Note that the preceding proof, particularly the sufficient condition, applies only when $H$ contains hyperedges of size at most $3$. Consider a size-$4$ hyperedge, $e=tuvw$, for example. Lifting the central vertex “$\times$” in its factor graph would create two edges, $tv$ and $uw$ w.l.o.g. However, two distinct edges cannot be derived from a single hyperedge $e$ through coalescence and dewetting, and this would invalidate the sufficient condition in Theorem II.2. This suggests that the connection between hypergraph and ordinary graph immersion is indeed more subtle, and one cannot simply reduce a hypergraph immersion problem to ordinary graph immersion for the general case. Therefore, alternative techniques are required.
Section III More Examples on $K_{4}^{2}$ and $K_{4}^{3}$
In this section, we give more nontrivial examples when the hypergraph has an immersion of $K_{4}^{2}$ and/or $K_{4}^{3}$.
Figure S3: There exist both immersions of $K_{4}^{2}$ and $K_{4}^{3}$ in an infinite honeycomb hyper-lattice, but of different depth.
Now, if we consider the honeycomb lattice again, but instead of pairwise connections, vertices are connected by hyperedges of size $6$ (or hexagons), we can show that an immersion of $K_{4}^{3}$ will occur; see Fig. S3 for the construction. We denote such structure as infinite honeycomb hyper-lattice. Following a similar idea to the proof of Theorem .2 in the main text, we can show that an immersion of $K_{4}^{2}$ still exists; see Fig. S3 for the construction. It is clear that for every image of $K_{4}^{2}$ in an infinite honeycomb hyper-lattice, there is an immersion of $K_{4}^{3}$, but it is not necessarily true for the opposite; see Fig. S3 for an example when three out of four images of the four vertices in $K_{4}^{2}$ are (directly) connected with the remaining one. Specifically, if we restrict the honeycomb hyper-lattice to a smaller region where we cannot find four vertices such that (i) none of them are on the boundary, (ii) any two of them do not share a hexagon, and (iii) at most two of three hexagons that are incident on three of the four vertices are on the boundary, then we cannot find an immersion of $K_{4}^{2}$. Hence, if we consider hypergraphs that are locally honeycomb hyper-lattice like (satisfying the above condition), and globally tree like, then an immersion of $K_{4}^{2}$ does not exist but $K_{4}^{3}$.
We finish this section with a slightly different type of graphs consisting of hyperedges of size $4$ that is necessarily shown in 3-dimensional space (or is not planar as before and in the main text).
Figure S4: Cuboid full lattice versus cuboid alternating lattice. Hypergraphs in the shape as a cuboid where there is no edges inside but on the four outer faces: (top) four faces are filled with hyperedges of size $4$, cuboid full lattice; (bottom) there can only be on hyperedge in the two opposite faces, cuboid alternating lattice.
Specifically, we consider the hypergraph in the shape as an infinite cuboid where there is no edges inside but on the four infinite faces; see Fig. S4 for part of it. We first consider the case when the four faces are filled with hyperedges (i.e., the top of Fig. S4), and we denote such structure as infinite cuboid full lattice hereafter. There is an immersion of $K_{4}^{2}$ in an infinite cuboid full lattice, since there exists an ordinary subgraph obtained with a series of dewetting operations on the hypergraph that is not series-parallel. Hence, there is an embedding of $K_{4}^{2}$, thus an immersion of $K_{4}^{2}$, in such hypergraph.
However, we now consider the case when the hyperedges only occur in an alternating manner such that for the two opposite faces, there can only be one hyperedge in the two faces (i.e., the bottom of Fig. S4), and we denote such structure as infinite cuboid alternating lattice hereafter. Then, we cannot find an immersion of $K_{4}^{2}$.
When it comes to $K_{4}^{3}$, it has an immersion in both hypergraphs; see the bottom of Fig. S4 for an example where vertices $A,B,C,D$ correspond to the four vertices in $H=K_{4}^{3}$. We also note that the infinite cuboid alternating lattice is a subgraph of the infinite cuboid full lattice.